# WonderHuman: Hallucinating Unseen Parts in Dynamic 3D Human Reconstruction

### v1

@@ Abstract
Monocular Video, 3D Gaussian Splatting, Human Unseen Part Reconstruction, Diffusion, Score Distillation Sampling.
@@ 1 Introduction
Existing avatar reconstruction methods have difficulty in reconstructing unseen parts of the human body. Previous methods [1, 2, 3, 4] typically rely on dense, synchronized multi-view inputs for the avatar reconstruction task. Recent advancements in implicit neural radians fields [5, 1, 3, 4] and 3D Gaussian Splatting [6, 7, 8, 9] have explored the high-fidelity reconstruction of both geometry and appearance of dynamic human bodies from relatively sparse multi-view videos. To reconstruct from monocular videos, other recent methods [10, 11, 12, 13, 14, 15, 16] reconstruct dynamic avatars by animating them within a canonical space derived from observation spaces using video frames. These works enable learning the inter-frame deformation to reconstruct a completed human avatar from the monocular videos. However, these methods still require the video to have full-view coverage of the human body, and typically fail to reconstruct unseen parts in the monocular video. Unfortunately, one often only has access to partial-view videos with limited viewpoints, such as front-view videos, leaving most parts of the human body unseen. Reconstructing these occluded parts thus poses a significant challenge for current methodologies.
@@
To address this challenge, we introduce WonderHuman to achieve high-quality avatar reconstruction from partial-view monocular videos. The key idea of WonderHuman is to hallucinate the unseen parts of the human using the generative prior encoded by large-scale image diffusion models such as Zero123 [17]. The hallucinations are then combined with a Gaussian Splatting [6]-based dynamic human reconstruction framework to get a full-body avatar.
@@
However, combining diffusion-generative priors in dynamic human reconstruction is not a trivial task with two outstanding challenges. First, the existing image diffusion generative models are designed mainly to produce single-view static images. Thus, maintaining visually accurate generated content and consistency across frames for dynamic human bodies using these generative priors is challenging; For instance, unrealistic artifacts such as blurs often appear when animating the generated bodies. Some existing works [18, 19, 20] can produce human bodies from single-view images using diffusion models, but they fail to handle dynamic cases (See Appendix B.1 for more details). Second, it’s challenging to ensure that the occluded or invisible portions of the human body generated by diffusion models are consistent with the observed visible parts. Any inconsistency between these generated and visible segments can significantly deteriorate the rendering quality of the human avatar, leading to visually incoherent results.
@@
To tackle these issues, we present WonderHuman for high-quality dynamic human reconstruction from monocular videos. We propose leveraging the generative priors embedded in a 2D diffusion model, trained on condensed images, to infer the unseen parts of the 3D human through distillation during reconstruction. We further introduce a novel Dual-Space Optimization method to ensure visual plausibility and consistency for dynamic human representations. Our Dual-Space Optimization utilizes Score Distillation Sampling (SDS) [21] in both the canonical and observation spaces. This approach ensures that the generated content remains natural and complete by accounting not only for the information in a canonical pose but also for the dynamics across poses in the observation space. This significantly enhances the rendering quality when animating the reconstructed human avatar. Moreover, a view selection strategy and a pose feature injection approach are employed to reconcile conflicts between the SDS predictions and the given information and fuse pose-dependent effects, enhancing dynamic synthesis and overall avatar fidelity.
@@
We conduct extensive experiments to validate the effectiveness of our method across broad benchmarks including ZJU-Mocap dataset [1], Monocap dataset [2], MVHumanNet [22] and In-the-wild dataset [23]. Compared to state-of-the-art methods [10, 15, 24, 25, 26], WonderHuman produces higher-quality photorealistic renderings of reconstructed human avatars, particularly in rendering visually plausible content for previously unseen parts of the human body. To summarize, our contributions are as follows:
@@ 2.1 Video-based Human Avatar Reconstruction
Recently, video-based avatar reconstruction methods primarily rely on regression-based approaches [27, 28, 29, 30, 31, 32, 33, 34] or the explicit tracking of human bodies [35, 36, 37, 38, 39, 40, 41]. Since the prosperity of Neural Rendering [5], many works [42, 43, 44, 1, 2, 4, 45, 46, 47, 11, 3, 12, 10, 48, 13] try to combine neural representations with human reconstructions. These methods associate implicit neural fields on human templates like SMPL [49]. While neural representations have strong representation ability, they are slow in training. But other works [50, 14, 51, 15] additionally introduced explicit representations like meshes [50, 14], and points [51] to improve its efficiency. Yet, achieving high-quality reconstruction results using neural radiance fields still requires neural networks that are expensive to train and render. Recently, Gaussian Splatting [6] has emerged as a prominent technology, as it efficiently represents and renders complex scenes with reduced training time, without compromising quality for speed. Many recent works [7, 52, 8, 9, 53, 54, 55, 16, 56, 57, 58, 59, 60, 61, 62] try to combine Gaussian Splatting in the avatar reconstruction, which allows efficient avatar rendering in real-time. In this paper, we focus on reconstructing human avatars from monocular video. In GaussianAvatar [16], 3D Gaussians are integrated with SMPL [49] to explicitly represent humans in various poses and clothing styles. SplattingAvatar [24] embeds Gaussians onto human triangle meshes, forming a hybrid representation that significantly enhances rendering speed. Furthermore, ExAvatar [25] extends this representation to reconstruct animatable hand poses and facial expressions. However, those methods require the input video to have full-view coverage of the human body and failed to generate unseen parts in the monocular video.
@@ 3 Preliminaries
where $I$ is the rendered image, $r$ is a set of query rays crossing the scene, and $Splatting(\cdot)$ is a differentiable rendering process. We refer readers to Kerbl et al.’s paper [6] for the details of Gaussian splatting.
@@ 4.2 Stage II: Invisible Appearance Reconstruction
Stage I produces an animatable 3D human model with visible appearances learned from partial-view video data, but the unobserved regions of the body typically suffer from relatively low visual quality. To ensure multi-view consistency for the unseen parts, we introduce a viewpoint-conditioned diffusion model as supervision, leveraging generative priors to predict the unseen views from the given inputs. Subsequently, we optimize the Gaussian decoder $G_{\theta}$ to reconstruct a fully renderable 3D human model from any viewpoint. To effectively utilize the observations and improve consistency between observed and hallucinated results, we introduce Dual-space Optimization, View Selection and Pose Feature Injection techniques in the following.
Figure 2: Left side: Dual-pace Optimization (a) w/o dual space optimization; (b) w/ canonical optimization only; (c) w/ dual-space optimization; Right side: Pose Feature Injection (d) ground truth; (e) w/o pose feature injection; (f) w/ pose feature injection.
@@ 4.2.1 Dual-space Optimization
Zero123 [17], a viewpoint-conditioned diffusion model, is used to hallucinate full-body views from partial video frames, using reference frames from the monocular video and target view camera parameters as conditioning inputs. Its explicit view control enables precise multi-view predictions for 3D reconstruction. We leverage Score Distillation Sampling (SDS) [21] loss for predicting the unseen parts of our 3D Gaussian human model in the observation space. Unfortunately, naively combining Zero123 using SDS for dynamic human reconstruction leads to unrealistic reconstruction results. For instance, directly applying SDS in canonical space often results in degenerated quality issues in avatars—when generating 3D models with 2D diffusion models (See Fig. 2(b)).
To address this, we introduce Dual-Space Optimization, which performs SDS optimization in both canonical and observation spaces. When conducting optimization in the canonical space, we use the rendering in the canonical space from Stage I as a conditioning reference for the 2D generative diffusion model. When conducting optimization in the observation space, we utilize the selected input images from the partial-view video as conditioning references.
The SDS optimization process, combining Zero123 with the dual-space strategy, is thus expressed as:
@@
where $I_{\theta}$ represents a generated image from an unseen view in observation or canonical space. $\epsilon_{\phi}$ is the predicted noise by Zero123 conditioned on the image $y$ and the target view camera parameters $(R,T)$.
Since $\mathcal{L}_{SDS}$ is applied in both canonical and observation spaces, we take the observed frames from the input video as $y_{image}$ when optimizing in observation space, and take the canonical rendering from Stage I as $y_{image}$ when optimizing in canonical space. This approach allows us to more effectively associate features across frames for the reconstruction of unseen parts.
During dual-space optimization, we found that appropriately balancing the training processes of the two spaces improves performance. As mentioned earlier, diffusion models face degeneration issues when optimizing in the canonical space. For instance, they struggle to predict accurate appearances for complex human poses in the observation space, often producing unrealistic ’tattoo-like’ appearances, as shown in Fig. 6(c). To address this, we set the weight between canonical and observation optimization as a hyperparameter in Stage II to enhances the overall process performance. This refined balance ensures better alignment of the model with the desired objectives, leading to more accurate and reliable outcomes in Stage II.
@@ 4.2.2 View Selection
In both canonical and observation space optimizations, we identify the invisible views that require refinement. By utilizing the differentiable rasterization of Gaussian Splatting [6], we determine the first intersecting Gaussian for each ray, marking these as visible points. Subsequently, visibility maps are rendered to differentiate between the visible and invisible regions of the human avatar as defined in Stage I. Specificly, we first estimate the visibility of each Gaussain. During the training of Stage I with seen views, given a ray $r$, the first Gaussian hit by the ray, $x$, is marked as a seen Gaussian, and its visibility $\psi$ is set to 1. Formally:
@@
where $\mathbb{I}_{v}=0$ signifies an unseen viewpoint, indicating that the avatar needs to be refined from viewpoint $v$ in Stage II.
@@ 4.3 Training Losses
In Stage I, we are modeling a dynamic avatar from partial-view videos using a Gaussian decoder. Additionally, we refine the input pose to correct inaccuracies from SMPL fitting. This stage utilizes MSE loss, SSIM loss [93], and perceptual LPIPS loss [94] between the predicted RGB images and ground truth, as $\mathcal{L}_{rgb}$, $\mathcal{L}_{ssim}$, and $\mathcal{L}_{lpips}$, respectively. We also apply Frobenius Norm loss as regularization terms for optimizable canonical features $S$, offset $\Delta x$, and scale $s$:
@@
where $\mathcal{L}_{SDS}^{o}$ and $\mathcal{L}_{SDS}^{c}$ represent the SDS loss for observation space and canonical space, respectively, as defined in Eq. (9).
Furthermore, we design a progressive training strategy in this stage, gradually diminishing the weight of SDS loss. This strategy is employed to enhance further the effectiveness and efficiency of the visible appearance reconstruction. More details on progressive training are in Appendix A.2.2.
@@
ZJU-Mocap(revised) MVhumanNets Monocap
[…]
HumanNeRF 19.21 0.9456 0.0715 19.38 0.9416 0.0706
Instant-NVR 19.90 0.9458 0.0630
SplattingAvatar 19.37 0.9436 0.0702 19.49 0.9467 0.0689 19.51 0.9444 0.0697
EXAvatar 19.65 0.9449 0.0678 19.69 0.9470 0.0671 19.65 0.9475 0.0676
GaussianAvatar 19.50 0.9434 0.0687 19.70 0.9471 0.0645 19.67 0.9489 0.0635
GuessTheUnseen 20.06 0.9493 0.0615 20.56 0.9502 0.0598
Ours 20.82 0.9552 0.0569 20.98 0.9517 0.0553 21.16 0.9532 0.0549
@@ 5.1 Datasets
ZJU-Mocap(revised) dataset [1]. This dataset is a multi-view dataset. We train and test using this dataset following Instant-NVR [15]. One specific camera capture is used as monocular training input and six cameras, evenly distributed around the object, are reserved for a comprehensive evaluation.
@@ 5.2 Implementation Details
All videos from all datasets are clipped to 3-5 seconds (100-150 frames) and exclusively capture front views of the subjects. During stage I, training is conducted on a single RTX-3090 GPU with a batch size of 2, requiring approximately 1 hour for 200 training epochs. In Stage II, the entire framework is trained on two RTX-3090 GPUs with a batch size of 1, while the diffusion model is loaded exclusively on the second GPU. The training process requires approximately 2–3 hours for 400 epochs, depending on the resolution.
@@ 5.3 Comparisons with Video-based Methods
[…]
HumanNeRF 30.23 0.9756 0.0314
SplattingAvatar 28.28 0.9693 0.0286
ExAvatar 29.46 0.9709 0.0253
GaussianAvatar 29.96 0.9716 0.0220
GuessTheUnseen 29.35 0.9685 0.0256
Ours 29.76 0.9712 0.0222
TABLE II: Quantitative evaluation on In-the-wild datasets (seen view only)

### v2

@@ Abstract
Monocular Video, Monocular Reconstruction, Gaussian Splatting, Human Reconstruction, Diffusion, Score Distillation Sampling.
@@ 1 Introduction
Existing avatar reconstruction methods have difficulty in reconstructing unseen parts of the human body. Previous methods [1, 2, 3, 4] typically rely on dense, synchronized multi-view inputs for the avatar reconstruction task. Recent advancements in implicit neural radiance fields [5, 1, 3, 4] and 3D Gaussian Splatting [6, 7, 8, 9] have explored the high-fidelity reconstruction of both geometry and appearance of dynamic human bodies from relatively sparse multi-view videos. To reconstruct from monocular videos, other recent methods [10, 11, 12, 13, 14, 15, 16] reconstruct dynamic avatars by animating them within a canonical space derived from observation spaces using video frames. These works enable learning the inter-frame deformation to reconstruct a completed human avatar from the monocular videos. However, these methods still require the video to have full-view coverage of the human body, and typically fail to reconstruct unseen parts in the monocular video. Unfortunately, one often only has access to partial-view videos with limited viewpoints, such as front-view videos, leaving most parts of the human body unseen. Reconstructing these occluded parts thus poses a significant challenge for current methodologies.
To address this challenge, we introduce WonderHuman to achieve high-quality avatar reconstruction from partial-view monocular videos.
@@
The key idea of WonderHuman is to hallucinate the unseen parts of the human using the generative prior encoded by large-scale image diffusion models such as Zero123 [17]. The hallucinations are then combined with a Gaussian Splatting [6]-based dynamic human reconstruction framework to obtain a full-body avatar.
@@
However, combining diffusion-based generative priors in dynamic human reconstruction is not a trivial task with two outstanding challenges. First, the existing image diffusion generative models are designed mainly to produce single-view static images. Thus, maintaining visually accurate generated content and consistency across frames for dynamic human bodies using these generative priors is challenging. For instance, unrealistic artifacts such as blurs often appear when animating the generated bodies. Some existing works [18, 19, 20] can produce human bodies from single-view images using diffusion models, but they fail to handle dynamic cases (See Appendix B.1 for more details). Second, it is challenging to ensure that the occluded or invisible portions of the human body generated by diffusion models are consistent with the observed visible parts. Any inconsistency between these generated and visible segments can significantly deteriorate the rendering quality of the human avatar, leading to visually incoherent results.
@@
To tackle these issues, we present WonderHuman for high-quality dynamic human reconstruction from monocular videos. We propose leveraging the generative priors embedded in a 2D diffusion model, trained on condensed images, to infer the unseen parts of the 3D human through distillation during reconstruction. We further introduce a novel Dual-space Optimization method to ensure visual plausibility and consistency for dynamic human representations. Our Dual-space Optimization utilizes Score Distillation Sampling (SDS) [21] in both the canonical and observation spaces. This approach is designed to bridge the gap between observation and canonical spaces by effectively fusing detailed appearance information captured from multi-pose observations into a consistent canonical representation. This process not only preserves spatial fidelity but also enhances temporal coherence (frame-by-frame) by modeling dynamic changes over time, leading to more stable and accurate reconstructions of animatable human avatars.
Moreover, a view selection strategy and a pose feature injection approach are employed to reconcile conflicts between the SDS predictions and the given information and fuse pose-dependent effects, enhancing dynamic synthesis and overall avatar fidelity.
@@
We conduct extensive experiments to validate the effectiveness of our method across broad benchmarks including ZJU-Mocap dataset [1], Monocap dataset [2], MVHumanNet [22], and In-the-wild dataset [23]. Compared to state-of-the-art methods [10, 15, 24, 25, 26], WonderHuman produces higher-quality photorealistic renderings of reconstructed human avatars, particularly in rendering visually plausible content for previously unseen parts of the human body. To summarize, our contributions are as follows:
@@ 2.1 Video-based Human Avatar Reconstruction
Recently, video-based avatar reconstruction methods primarily rely on regression-based approaches [27, 28, 29, 30, 31, 32, 33, 34] or the explicit tracking of human bodies [35, 36, 37, 38, 39, 40, 41]. Since the prosperity of Neural Rendering [5], many works [42, 43, 44, 1, 2, 4, 45, 46, 47, 11, 3, 12, 10, 48, 13] try to combine neural representations with human reconstructions. These methods associate implicit neural fields with human templates such as Skinned Multi-Person Linear model (SMPL) [49], a widely used parametric 3D human body model that represents pose and shape using a learned mesh deformation framework. While neural representations have strong representation ability, they are slow in training. But other works [50, 14, 51, 15] additionally introduced explicit representations like meshes [50, 14], and points [51] to improve its efficiency. Yet, achieving high-quality reconstruction results using neural radiance fields still requires neural networks that are expensive to train and render. Recently, Gaussian Splatting [6] has emerged as a prominent technology, as it efficiently represents and renders complex scenes with reduced training time, without compromising quality for speed. Many recent works [7, 52, 8, 53, 54, 55, 16, 56, 9, 57, 58, 59, 60] try to combine Gaussian Splatting in the avatar reconstruction, which allows efficient avatar rendering in real-time. In this paper, we focus on reconstructing human avatars from monocular video. In GaussianAvatar [16], 3D Gaussians are integrated with SMPL [49] to explicitly represent humans in various poses and clothing styles. SplattingAvatar [24] embeds Gaussians onto human triangle meshes, forming a hybrid representation that significantly enhances rendering speed. Furthermore, ExAvatar [25] extends this representation to reconstruct animatable hand poses and facial expressions. Building on prior advances, GPS-Gaussians+ [61] extend the capability of GPS-Gaussian [9] from dense-view input to sparse-view input. TaoAvatar [62] delivers real-time, high-fidelity full-body avatars capable of running on mobile and wearable devices, bridging the gap between quality and efficiency. And MIMO [63] achieves avatar generation, motion generation, and user interaction in a unified framework. Vid2Avatar [64] addresses overfitting issues by using a universal model for avatar reconstruction from monocular video. However, those methods require the input video to have full-view coverage of the human body and fail to generate unseen parts in the monocular video.
@@ 3 Preliminaries
where II is the rendered image, $r$ is a set of query rays crossing the scene, and $Splatting(\cdot)$ is a differentiable rendering process. We refer readers to Kerbl et al.’s paper [6] for the details of Gaussian splatting.
@@ 4.2 Stage II: Invisible Appearance Reconstruction
Stage I produces an animatable 3D human model with visible appearances learned from partial-view video data, but the unobserved regions of the body typically suffer from relatively low visual quality. To ensure multi-view consistency for the unseen parts, we introduce a viewpoint-conditioned diffusion model as supervision, leveraging generative priors to predict the unseen views from the given inputs. Subsequently, we optimize the Gaussian decoder $G_{\theta}$ to reconstruct a fully renderable 3D human model from any viewpoint. To effectively utilize the observations and improve consistency between observed and hallucinated results, we introduce Dual-space Optimization, View Selection, and Pose Feature Injection techniques in the following.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 2: Left side: Dual-space Optimization (a) w/o Dual-space Optimization; (b) w/ canonical optimization only; (c) w/ Dual-space Optimization; Right side: Pose Feature Injection (d) ground truth; (e) w/o pose feature injection; (f) w/ pose feature injection.
@@ 4.2.1 Dual-space Optimization
Zero-1-to-3 [17], a viewpoint-conditioned diffusion model, is used to hallucinate full-body views from partial video frames, using reference frames from the monocular video and target view camera parameters as conditioning inputs. Its explicit view control enables precise multi-view predictions for 3D reconstruction. We leverage Score Distillation Sampling (SDS) [21] loss for predicting the unseen parts of our 3D Gaussian human model in the observation space. Unfortunately, naively combining Zero123 using SDS for dynamic human reconstruction leads to unrealistic reconstruction results. For instance, directly applying SDS in canonical space often results in quality degeneration in avatars—when generating 3D models with 2D diffusion models (See Fig. 2(b)).
To address this, we introduce Dual-space Optimization, which performs SDS optimization in both canonical and observation spaces. When conducting optimization in the canonical space, we use the rendering in the canonical space from Stage I as a conditioning reference for the 2D generative diffusion model. When conducting optimization in the observation space, we utilize the selected input images from the partial-view video as conditioning references. The SDS optimization process, combining Zero-1-to-3 with the Dual-space strategy, is thus expressed as:
@@
where $I_{\theta}$ represents a generated image from an unseen view in observation or canonical space. $\epsilon_{\phi}$ is the predicted noise by Zero123 conditioned on the image $y$ and the target view camera parameters $(R,T)$. Since $\mathcal{L}_{SDS}$ is applied in both canonical and observation spaces, we take the observed frames from the input video as $y_{image}$ when optimizing in observation space, and take the canonical rendering from Stage I as $y_{image}$ when optimizing in canonical space. This approach allows us to more effectively associate features across frames for the reconstruction of unseen parts.
During Dual-space Optimization, we found that appropriately balancing the training processes of the two spaces improves performance. As mentioned earlier, diffusion models face degeneration issues when optimizing in the canonical space. For instance, they struggle to predict accurate appearances for complex human poses in the observation space, often producing unrealistic ’tattoo-like’ appearances, as shown in Fig. 9(c). To address this, we set the weight between canonical and observation optimization as a hyperparameter in Stage II to enhance the overall process performance. This refined balance ensures better alignment of the model with the desired objectives, leading to more accurate and reliable outcomes in Stage II.
@@ 4.2.2 View Selection
In both canonical and observation space optimizations, we identify the invisible views that require refinement. By utilizing the differentiable rasterization of Gaussian Splatting [6], we determine the first intersecting Gaussian for each ray, marking these as visible points. Subsequently, visibility maps are rendered to differentiate between the visible and invisible regions of the human avatar as defined in Stage I. Specificly, we first estimate the visibility of each Gaussian. During the training of Stage I with seen views, given a ray $r$, the first Gaussian hit by the ray, $x$, is marked as a seen Gaussian, and its visibility $\psi$ is set to 1. Formally:
@@
where $\mathbb{I}_{v}=0$ signifies an unseen viewpoint, indicating that the avatar needs to be refined from viewpoint $v$ in Stage II. And visibility map examples are shown in Fig. 3.
((a))
((b))
@@ 4.3 Training Losses
In Stage I, we are modeling a dynamic avatar from partial-view videos using a Gaussian decoder. Additionally, we refine the input pose to correct inaccuracies from SMPL fitting. This stage utilizes mean squared error(MSE) loss, structural similarity(SSIM) loss [95], and perceptual similarity(LPIPS) loss [96] between the predicted RGB images and ground truth, as $\mathcal{L}_{rgb}$, $\mathcal{L}_{ssim}$, and $\mathcal{L}_{lpips}$, respectively. We also apply Frobenius Norm loss as regularization term for optimizable canonical features $S$, offset $\Delta x$, and scale $s$:
@@
where $\mathcal{L}_{SDS}^{o}$ and $\mathcal{L}_{SDS}^{c}$ denote the SDS loss in observation and canonical space, respectively, as defined in Eq. (9).
Furthermore, we design a progressive training strategy in this stage, gradually diminishing the weight of SDS loss. This strategy is employed to further enhance the effectiveness and efficiency of the visible appearance reconstruction. More details on progressive training are in Appendix A.2.2.
@@ 5.1 Datasets
ZJU-Mocap(revised) dataset [1]. This dataset is a multi-view dataset. We train and test using this dataset following Instant-NVR [15]. One specific camera view is used as the monocular training input, while six cameras evenly distributed around the object are reserved for comprehensive evaluation.
@@
ZJU-Mocap(revised) Monocap MVHumanNets
[…]
HumanNeRF 19.21 0.9456 0.0715 19.38 0.9416 0.0706
Instant-NVR 19.90 0.9458 0.0630 20.03 0.9469 0.0613
SplattingAvatar 19.37 0.9436 0.0702 19.51 0.9444 0.0697 19.49 0.9467 0.0689
ExAvatar 19.65 0.9449 0.0678 19.65 0.9475 0.0676 19.69 0.9470 0.0671
GaussianAvatar 19.50 0.9434 0.0687 19.67 0.9489 0.0635 19.70 0.9471 0.0645
GuessTheUnseen 20.06 0.9493 0.0615 20.56 0.9502 0.0598
Ours 20.82 0.9552 0.0569 21.16 0.9532 0.0549 20.98 0.9517 0.0553
@@ 5.2 Implementation Details
All videos from all datasets are clipped to 3-5 seconds (100-150 frames) and exclusively capture front views of the subjects. During stage I, training is conducted on a single RTX-3090 GPU with a batch size of 2, requiring approximately 1 hour for 200 training epochs. In Stage II, the entire framework is trained on two RTX-3090 GPUs with a batch size of 1, while the diffusion model is loaded exclusively on the second GPU. Overall, the Gaussian parameter decoder consists of approximately 186K parameters. The training process requires approximately 28 GB of GPU memory. Training requires approximately 2–3 hours for 400 epochs, while inference operates at a speed of 13–18 frames per second, depending on the resolution.